OpenStudy (anonymous):

please help! Beth is two years older than Julio. Gerald is twice as old as Beth. Debra is twice as old as Gerald. The sum of their ages is 38. How old is Beth?

OpenStudy (anonymous):

B = J+2 G = 2B D= 2G B+J+D+G = 38 There are your equations. Solve for B.